The sum of the digits of a two-digital numeral is 11. The number is 9 less than the number named when the digits are reversed. Find the numerals.
Equation & answer please! Thanks!

Answer by ankor@dixie-net.com(22740)

let a = 10's digit
let b = the units digit
then
10a+b = the number
:
Write an equation for each statement
:
The sum of the digits of a two-digital numeral is 11.
a + b = 11
:
The number is 9 less than the number named when the digits are reversed.
10a + b = 10b + a - 9
10a - a + b - 10b = -9
9a - 9b = -9
simplify divide by 9, then add the 1st equation
a - b = -1
a + b = 11
-------------adding eliminates b, find a
2a + 0 = 10
a = 10/2
a = 5
I'll let you find b, check your answer in the 2nd statement